At what angle to the vertical must a motorcyclist lean so as to negotiate a bend with a radius of 280m at a speed of 180km/h.


1
Expert's answer
2021-07-21T09:06:21-0400

Let's apply Newton's Second Law of Motion in projections on axis "x"- and "y":


"Ncos\theta=mg,""Nsin\theta=\dfrac{mv^2}{r}."

Dividing the second equation by the first one, we get:


"tan\theta=\dfrac{v^2}{rg},""\theta=tan^{-1}(\dfrac{v^2}{rg}),""\theta=tan^{-1}(\dfrac{(50\ \dfrac{m}{s})^2}{280\ m\cdot9.8\ \dfrac{m}{s^2}})=42^{\circ}."